Calculating equilibrium concentrations when the net reaction proceeds in reverse Consider mixture C, which will cause the net reaction to proceed in reverse. Concentration (M)initial:change:equilibrium:[XY]0.200+x0.200+x←net⇌[X]0.300−x0.300−x+[Y]0.300−x0.300−x The change in concentration, x, is positive for the reactants because they are produced and negative for the products because they are consumed. Part C Based on a Kc value of 0.160 and the data table given, what are the equilibrium concentrations of XY, X, and Y, respectively? Express the molar concentrations numerically.
Answer #1

XY --------> X + Y

Kc = [X] [Y] / [XY]

0.16 = [0.3-x] [ 0.3-x] / [0.2+x]

0.032 + 0.16x = 0.09 + x^2 - 0.6x

x^2 - 0.76x + 0.058 = 0

x = 0.086

COncentration of XY = 0.2+x = 0.286

Concentration of X = 0.3-x = 0.214

Concentration of Y = 0.3-x = 0.214